# Quantum Chemical and Trajectory Surface Hopping Molecular Dynamics Study of Iodine-based BODIPY Photosensitizer

### v1

@@ 3.6 Triplet quantum yield
Triplet quantum yield is a useful concept to quantitatively assess the efficiency of triplet states generation in photophysical processes. 111, 112 It can be expressed as
@@ 4 Conclusions
A computational study of I-BODIPY (2-ethyl-4,4-difluoro-6,7-diiodo-1,3-dimethyl-4-bora-3a,4a-diaza-s-indacene) was carried out to investigate its key photophysical properties as a potential triplet photosensitizer capable of generating singlet oxygen. Multireference CASSCF, CASPT2 and DMRG methods were used to verify and prove the applicability of the single-reference ADC(2) method for studying the electronic structure of the excited states of the molecule, at least in the vicinity of its ground- and (lowest-lying) excited-state equilibrium geometries. Careful benchmarking of several exchange-correlation functionals in the TD-DFT/TDA framework was done with respect to ADC(2), CASPT2 and CASSCF with a particular attention paid to the correct description of the excited states dominated by charge-transfer excitations as well as to the ability of TD-DFT employing a basis set with a two-component pseudopotential on the iodine atoms to well capture SOCs between the low-lying singlets and triplets. The magnitudes of SOC between excited electronic states of all types found were thoroughly discussed using the Slater–Condon rules applied to an arbitrary one-electron one-center effective spin–orbit Hamiltonian. The geometry dependence of SOCs between the lowest-lying states was also addressed. Based on these investigations, the TD-DFT/B3LYP and TD-DFT(TDA)/BHLYP approaches were selected as the methods of choice for the subsequent absorption spectra and TSH MD simulations, respectively.

### v2

@@ 3.6 Triplet quantum yield
Triplet quantum yield is a useful concept to quantitatively assess the efficiency of triplet states generation in photophysical processes. 117, 118 It can be expressed as
@@ 4 Conclusions
A computational study of I-BODIPY (2-ethyl-4,4-difluoro-6,7-diiodo-1,3-dimethyl-4-bora-3a,4a-diaza-s-indacene) was carried out to investigate its key photophysical properties as a potential triplet photosensitizer capable of generating singlet oxygen. Multireference CASSCF, CASPT2 and DMRG methods were used to verify and prove the applicability of the single-reference ADC(2) method for studying the electronic structure of the excited states of the molecule, at least in the vicinity of its ground- and (lowest-lying) excited-state equilibrium geometries. Careful benchmarking of several exchange-correlation functionals in the TD-DFT framework was done with respect to ADC(2), CASPT2 and CASSCF with a particular attention paid to the correct description of the excited states dominated by charge-transfer excitations as well as to the ability of TD-DFT employing a basis set with a two-component pseudopotential on the iodine atoms to well capture SOCs between the low-lying singlets and triplets. The magnitudes of SOC between excited electronic states of all types found were thoroughly discussed using the Slater–Condon rules applied to an arbitrary one-electron one-center effective spin–orbit Hamiltonian. The geometry dependence of SOCs between the lowest-lying states was also addressed. Based on these investigations, the TD-DFT/B3LYP and TD-DFT(TDA)/BHLYP approaches were selected as the methods of choice for the subsequent NEA absorption spectra simulations and mixed quantum-classical TSH MD simulations, respectively.